### v2

Abstract
We provide a framework to study stability notions for two-sided dynamic matching markets in which matching is one-to-one and irreversible. The framework gives centerstage to the set of matchings an agent anticipates would ensue should they remain unmatched, which we refer to as the agent’s conjectures. A collection of conjectures, together with a pairwise stability and individual rationality requirement given the conjectures, defines a solution concept for the economy. We identify a sufficient condition–consistency–for a family of conjectures to lead to a nonempty solution (cf. Hafalir, 2008). As an application, we introduce two families of consistent conjectures and their corresponding solution concepts: continuation-value-respecting dynamic stability, and the extension to dynamic markets of the solution concept in Hafalir (2008), sophisticated dynamic stability.
JEL classification codes: D47, C78
Keywords: dynamic matching, matching with externalities, conjectures, dynamic stability
1 Introduction
In this paper, we provide a unifying framework to study solution concepts for two-sided, one-to-one, dynamic matching markets in which matching is irreversible. A key ingredient in any such solution concept is what matchings an agent expects would ensue should they remain unmatched and wait for a better matching opportunity; we dub this set of matchings the agent’s conjectures. Following the literature on matching with externalities (Sasaki and Toda, 1996; Hafalir, 2008), we take conjectures as a primitive and define a solution concept given these conjectures. We provide a sufficient condition these conjectures must satisfy for the solution concept to be nonempty. Armed with this result, we propose two solution concepts for dynamic matching markets with nonempty solutions.
Formally, we study a two-sided, dynamic matching market, in which agents on both sides of the economy arrive over time. We assume arrivals are deterministic, that is, agents can perfectly foresee when each agent arrives in the economy.Footnote 2Footnote 2Footnote 2We assume deterministic arrivals because it simplifies notation, but it is not needed for our results. See 2.2. A matching specifies who matches with who in each period, subject to two constraints: agents cannot be matched before they arrive, and agents can have at most one matching partner on the opposite side. Agents are discounted utility maximizers, and their payoff from a matching solely depends on who they match with and when. Our goal is to study solution concepts for this economy.
In Section 3, we introduce the two main ingredients of the model: an agent’s conjectures (5) and a solution concept given the conjectures (6). When determining whether a matching is self-enforcing, and hence a solution for the economy, the matching should at the very least be immune to two kinds of blocks: blocks by individual agents who may not find their current matching partner acceptable, and blocks by pairs of contemporaneous agents who would rather match together. As we explain next, an agent’s conjectures allow us to define the agent’s set of acceptable matching partners. We say a matching is a solution for the economy given the conjectures if no agent is matched to an unacceptable matching partner, and no two agents who are available to match at the same time prefer to match together. As we explain in Section 3, in one-period economies, the solution concept given the conjectures reduces to the notion of stability in Gale and Shapley (1962), independently of the conjectures.
An agent’s conjectures describe the set of matchings the agent assumes may ensue should they remain unmatched in a given period. Importantly, a given conjectured matching describes both the current period’s matching outcome and the continuation matching from the next period onward. Because the set of agents available to match from the next period onward depends on both the newly arriving agents and those who remain unmatched from previous periods, the current period’s matching and the next period’s arrivals define the set of feasible continuation matchings. In other words, an agent’s outcome from remaining unmatched cannot be defined independently of other contemporaneous agents’ matching outcomes. Thus, having assumed away payoff externalities, an externality still exists in our model as other contemporaneous agents’ matching outcomes determine which continuation matchings are feasible. For a given conjectured matching, we can define an agent’s acceptable partners to be those preferred to the conjectured matching. Thus, to determine an agent’s acceptable partners, we need an assumption on how to select a continuation matching from the agent’s conjectures. Following Sasaki and Toda (1996), we say the agent finds a matching partner acceptable if the matching partner is preferred to some conjectured continuation matching.
The paper’s main result, Theorem 4.1, provides a sufficient condition on the family of conjectures for matchings satisfying the solution concept to exist. We refer to the conjectures satisfying this property as consistent. Consistent conjectures guarantee that a candidate matching, which can be computed from the model primitives, is a solution for the economy (see Section 4 for a definition of the set of candidate matchings). Loosely speaking, consistency requires the following condition holds between the candidate matching–denote it by $m^{\star}$–and the conjectures. In each period and for each agent who is unmatched under $m^{\star}$ in that period, the candidate matching $m^{\star}$ should be one of the agent’s conjectures. In other words, when checking the continuation candidate matching against their own conjectures for what matchings would ensue given that they remain unmatched, the agents who are unmatched under $m^{\star}$ will find the candidate matching to be consistent with their conjectures. Consistency is related to the condition in Hafalir (2008, Proposition 2), which ensures the existence of stable matchings in static markets with payoff externalities; we explain the connection between the two in Section 4.
Whereas consistency is defined relative to the candidate matching, an immediate corollary is that the following conjectures always lead to a nonempty solution concept. Suppose that each agent’s conjectures only rule out continuation matchings that do not conform with the solution concept in the continuation. For instance, in a two-period economy, these conjectures specify that in the last period only stable matchings among the remaining unmatched agents and the new arrivals are possible. These conjectures always satisfy consistency and Theorem 4.1 implies they lead to a non-empty solution concept. That is, if the agents conjecture that the economy conforms to the solution concept in the continuation and place no restrictions on the current period matching, the set of matchings that conform to the solution concept from today onward is nonempty.
Theorem 4.1 suggests a way to define solution concepts for dynamic matching markets with nonempty solution correspondences by focusing attention on conjectures that satisfy consistency. We provide two illustrations of this approach in Section 5, where we present two such conjectures and their corresponding solution concepts: continuation-value-respecting dynamic stability and sophisticated dynamic stability (Hafalir, 2008). Continuation-value-respecting dynamic stability (henceforth, CVR-dynamic stability) is a refinement of dynamic stability (Doval, 2022). Like the conjectures in the previous paragraph, the conjectures in CVR-dynamic stability are defined recursively with the solution concept, with agents ruling out continuation matchings that are not $\mathrm{CVR}$-dynamically stable. In contrast to the conjectures in the previous paragraph, CVR-dynamic stability also places restrictions on the current period matching. In particular, each agent rules out that other contemporaneous agents leave the economy matched to matching partners that are worse than their worst conjectured matching partner. That is, each agent’s conjectures respect other agents’ continuation values as specified by their own conjectures. Sophisticated dynamic stability extends the solution concept in Hafalir (2008), sophisticated expectations, to dynamic matching markets. Sophisticated dynamic stability builds on the rational expectations solution, a classical solution concept in matching with externalities that does not have consistent conjectures (see 4.1). Building on the insight of Hafalir (2008), the conjectures of sophisticated dynamic stability are iteratively built from those corresponding to rational expectations to ensure their consistency.
Related Literature
Our paper builds on the literature on two-sided matching with externalities (Li, 1993; Sasaki and Toda, 1996; Hafalir, 2008) to contribute to the literature on two-sided dynamic matching markets with irreversible matching (Doval, 2022; Nicolò et al., 2023, 2024).
We are closest to the contribution of Hafalir (2008). In a static, two-sided, one-to-one matching market with payoff externalities, Hafalir (2008) provides a sufficient condition on agents’ conjectures under which stable matchings exist. As we discuss in detail in Section 4, Theorem 4.1 provides the analogue of Hafalir (2008, Proposition 2) for dynamic matching markets. Whereas conceptual differences between the papers exist, the logic behind the constructions is the same. Furthermore, we provide the natural extension of the solution concept in Hafalir (2008), sophisticated expectations, to the dynamic environments under consideration in Section 5.2. We also introduce a new solution concept for dynamic matching markets, CVR-dynamic stability, which has no analogue in his work.
The literature on cooperative solution concepts for dynamic matching markets highlights the role of agents’ conjectures about what matching would ensue after a block in defining solution concepts (see, e.g., Kurino, 2009; Kadam and Kotowski, 2018; Kotowski, 2019; Altinok, 2019; Liu, 2023 for repeated matching markets and Doval, 2022; Nicolò, Salmaso, and Saulle, 2023, 2024 for irreversible matching, and Doval (Forthcoming) for a survey treatment of this literature). Because solution concepts in this literature are typically predicated on perfection requirements–i.e., the same solution concept applying at the appropriately defined continuation–agents’ conjectures are defined jointly with the solution concept. We follow a complementary approach in which we take agents’ conjectures as given and use them to define a solution concept. This allows us to identify a property of conjectures–consistency–that guarantees the corresponding solution concept is nonempty, and underlies existence results in the literature (cf. Doval, 2022). We also propose new solution concepts that refine those existing in the literature, illustrating how our result can be used as a recipe of sorts to define stability notions for dynamic matching markets.
Recently, Nicolò, Salmaso, and Saulle (2024) study a third refinement of dynamic stability (Doval, 2022). Like the conjectures in dynamic stability, the conjectures in the solution concept in Nicolò et al. (2024) are defined recursively based on what the solution concept specifies in the continuation. In particular, the agents rule out all matchings that do not abide by the solution concept from tomorrow onward. Unlike dynamic stability, the solution concept in Nicolò et al. (2024) imposes more structure on what matchings the agents can conjecture will ensue in the current period when considering remaining unmatched: The authors define conjectures in a given period through an iterative procedure which, at each stage, rules out current period matchings that do not satisfy some minimal stability requirements given the conjectures at that stage.Footnote 3Footnote 3Footnote 3Whereas the main focus of the paper is on the conjectures that obtain as the limit of this procedure, by stopping the procedure at a given stage the authors can study stability under different levels of sophistication (Hafalir, 2008). It is possible to show the conjectures they construct satisfy consistency. Following Doval (2022), their proof is based on showing that (the analogue for their solution concept of) the candidate matchings we define in Section 4 satisfy the conditions of their solution concept. Furthermore, the authors show via an example that the rational expectations solution may fail to exist in a two-sided dynamic matching market.
2 Model
For ease of comparison, we follow the notation in Doval (2022).
Two-sided economy
Throughout the paper, the focus is on two-sided economies, with sides $A$ and $B$, that last for $T<\infty$ periods. Denote by $\mathscr{A}$ the finite set of agents on side $A$, with elements $a\in\mathscr{A}$. Similarly, $\mathscr{B}$ denotes the finite set of agents on side $B$, with elements $b\in\mathscr{B}$. When we do not take a stance on an agent’s side, we denote by $k$ an element in $\mathscr{A}\cup\mathscr{B}$.
Arrivals
A length-$T$ economy, denoted $E^{T}=(A_{1},B_{1},\dots,A_{T},B_{T})$, consists of two sequences, $\{A_{1},\dots,A_{T}\}$ and $\{B_{1},\dots,B_{T}\}$, of disjoint subsets of $\mathscr{A}$ and $\mathscr{B}$, respectively.Footnote 4Footnote 4Footnote 4It is without loss to assume the sets $(A_{1},\dots,A_{T})$ are disjoint as we do not rule out indifferences. Thus, $a$ denotes a unique individual in the economy, but $a$ may be payoff-equivalent to other arriving agents on side $A$. The same applies for the sets $(B_{1},\dots,B_{T})$ being disjoint. For $t\in\{1,\dots,T\}$, $A_{t}$ are the side $A$-agents that arrive in period $t$, and $B_{t}$ are the side $B$-agents that arrive in period $t$. For any $t\leq T$, let $\overline{A}_{t}=\cup_{s=1}^{t}A_{s}$ denote the implied arrivals on side $A$ through period $t$; similarly, let $\overline{B}_{t}=\cup_{s=1}^{t}B_{s}$ denote the implied arrivals on side $B$ through period $t$.
Dynamic Matching
With this notation at hand, we define matchings in a given period (1) and matchings for the dynamic economy (2):
Definition 1 (Period-$t$ matching).
A period-$t$ matching is a mapping $m_{t}:\overline{A}_{t}\cup\overline{B}_{t}\mapsto\overline{A}_{t}\cup\overline{B}_{t}$ such that the following hold:
  1. 1.
    For all $a\in\overline{A}_{t}$, $m_{t}(a)\in\{a\}\cup\overline{B}_{t}$,
  2. 2.
    For all $b\in\overline{B}_{t}$, $m_{t}(b)\in\overline{A}_{t}\cup\{b\}$,
  3. 3.
    For all $k\in\overline{A}_{t}\cup\overline{B}_{t}$, $m_{t}(m_{t}(k))=k$.
Definition 2 (Irreversible dynamic matching).
A matching $m$ is a tuple $(m_{1},\dots,m_{T})$ such that
  1. 1.
    For all $t\in\{1,\dots,T\}$, $m_{t}$ is a period-$t$ matching,
  2. 2.
    For all $t\in\{1,\dots,T\}$, for all $a\in\overline{A}_{t}$, if $m_{t}(a)\neq a$, then $m_{s}(a)=m_{t}(a)$ for all $s\geq t$.
Let $\pazocal{M}_{T}$ denote the set of matchings.
Part 2 of the definition guarantees that once a pair $(a,b)$ is matched, they are matched in all subsequent periods. Given a matching $m\in\pazocal{M}_{T}$, let $m^{t-1}$ denote the tuple $(m_{1},\dots,m_{t-1})$, with $m^{0}=\{\emptyset\}$ and let $\pazocal{M}_{T}(m^{t-1})$ denote the set of matchings that coincide with $m$ through period $t-1$. For an agent $k\in\overline{A}_{t}\cup\overline{B}_{t}$, let $\pazocal{M}_{T}(m^{t-1},k)$ denote the subset of $\pazocal{M}_{T}(m^{t-1})$ such that $k$ is unmatched through period $t$.
Available agents and induced economies
Fix a matching $m\in\pazocal{M}_{T}$, and suppose agents have matched according to $m$ through period $t-1$. Then, the set of agents who can match in period $t$ is determined by the unmatched agents in period $t-1$ and the new arrivals in period $t$, $(A_{t},B_{t})$. Formally,
$\pazocal{A}_{t}(m^{t-1})$ =$=\{a\in\overline{A}_{t-1}:m_{t-1}(a)=a\}\cup A_{t},$
$\pazocal{B}_{t}(m^{t-1})$ =$=\{b\in\overline{B}_{t-1}:m_{t-1}(b)=b\}\cup B_{t}.$
For each $t\geq 1$, a matching $m$ induces a continuation economy of length $T-t$, $E_{t+1}^{T}(m^{t})=(\pazocal{A}_{t+1}(m^{t}),\pazocal{B}_{t+1}(m^{t}),\dots,A_{T},B_{T})$.
Preferences
We close the model by defining the agents’ preferences. Throughout, we assume agents discount the future. Formally, each $a\in\mathscr{A}$ defines a discount factor $\delta_{a}\in[0,1]$ and a Bernoulli utility, $u(a,\cdot):\mathscr{B}\cup\{a\}\mapsto\mathbb{R}$. Similarly, each $b\in\mathscr{B}$ defines a discount factor $\delta_{b}\in[0,1]$ and a Bernoulli utility, $v(\cdot,b):\mathscr{A}\cup\{b\}\mapsto\mathbb{R}$. We normalize to 00 the payoffs of remaining single for one period. That is, for all $a\in\mathscr{A}$ and all $b\in\mathscr{B}$, $u(a,a)=v(b,b)=0$.
Given a matching $m\in\pazocal{M}_{T}$, let $U_{t}(a,m)$ and $V_{t}(b,m)$ denote respectively $a$’s and $b$’s utilities of matching $m$ from period $t$ onwards. Formally, fix a matching $m\in\pazocal{M}_{T}$ and a period $t$. For every $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$, let $t_{m}(k)$ denote the first date at which $k$ is matched under $m$. That is, $t_{m}(k)$ is the smallest index $s$ such that $t\leq s$ and $m_{s}(k)\neq k$; otherwise, let $t_{m}(k)=T$. Then, for $a\in\pazocal{A}_{t}(m^{t-1})$, let
$U_{t}(a,m)=\delta_{a}^{t_{m}(a)-t}u(a,m_{T}(a)),$
denote $a$’s payoff from matching $m$ at date $t$. Similarly, for $b\in\pazocal{B}_{t}(m^{t-1})$, let
$V_{t}(b,m)=\delta_{b}^{t_{m}(b)-t}v(m_{T}(b),b),$
denote $b$’s payoff from matching $m$ at date $t$.
We record for future reference two properties a matching $m$ may satisfy:
Definition 3 (Individual rationality).
A matching $m\in\pazocal{M}_{T}$ is individually rational if for all $a\in\mathscr{A}$, $u(a,m_{T}(a))\geq 0$, and for all $b\in\mathscr{B}$, $v(m_{T}(b),b)\geq 0$.
Definition 4 (Static notion of stability).
Suppose $T=1$. A matching $m$ for economy $E^{1}=(A_{1},B_{1})$ is stable if the following hold:
  1. 1.
    $m$ is individually rational, and
  2. 2.
    no pair $(a,b)\in A_{1}\times B_{1}$ exists such that $u(a,b)>U_{1}(a,m)$ and $v(a,b)>V_{1}(b,m)$.
We denote by $\mathrm{S}(E^{1})$ the set of stable matchings for $E^{1}$.
Remark 2.1 (Notation).
We note that several pieces of notation, like $\pazocal{M}_{T}$, $\pazocal{A}_{t}(\cdot)$, $\pazocal{B}_{t}(\cdot)$, denote objects whose definitions depend on the economy $E^{T}$ under consideration. To be precise, we should note them $\pazocal{M}_{T}(E^{T}),\pazocal{A}_{t}(\cdot;E^{T}),\pazocal{B}_{t}(\cdot,E^{T})$. In what follows, we continue to omit this dependence whenever it does not lead to confusion.
Remark 2.2 (Simplifying assumptions).
We note two assumptions that simplify notation, but are otherwise not needed for the results. First, we assume agents can perfectly foresee who arrives in each period. Instead, our results extend to the case in which arrivals are stochastic so that an economy is defined as a distribution over sequences $E^{T}$ as in Doval (2021). Second, we assume time-discounted preferences, but all that matters for our results is that agents are dynamically consistent.
3 A solution concept given conjectures
In this section, we introduce the two key objects of our paper: an agent’s conjectures and a solution concept induced by these conjectures.
Fix an economy $E^{T}$, a matching $m$, a period $t$, and an agent $k$ who is available to match in period $t$. Agent $k$’s conjectures describe the matchings $k$ considers possible should they decide to remain unmatched in period $t$:
Definition 5 (Conjectures).
Fix a period $t$, a matching through $t-1$, $m^{t-1}$, and an agent $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$. A conjecture for agent $k$ is a subset $\varphi_{t}(m^{t-1},k)$ of $\pazocal{M}_{T}(m^{t-1},k)$.
Because the conjectures represent what $k$ expects to obtain should they remain unmatched in period $t$, each conjectured matching $\overline{m}$ satisfies $\overline{m}_{t}(k)=k$. Each conjectured matching $\overline{m}\in\varphi_{t}(m^{t-1},k)$ describes both the matching that ensues in period $t$, $\overline{m}_{t}$, and the continuation matching $(\overline{m}_{t+1},\dots,\overline{m}_{T})$. Whereas agent $k$’s matching outcome is determined by the continuation matching $(\overline{m}_{t+1},\dots,\overline{m}_{T})$, specifying who matches in period $t$ is necessary to define what matchings are feasible from period $t+1$ onward. For this reason, a conjecture describes the matching starting from period $t$ onward. That defining the set of matchings $k$ may face from period $t+1$ onward necessitates specifying the outcomes of the other period-$t$ agents implies an externality exists in this economy, even if no direct payoff externalities exist.
Given a family of conjectures, $\Phi\equiv\{\varphi_{t}(m^{t-1},k):1\leq t\leq T,m^{t-1},k\in\mathscr{A}\cup\mathscr{B}\}$, one can define a solution concept for $E^{T}$, as follows:
Definition 6 (Solutions induced by $\Phi$).
A matching $m\in\pazocal{M}_{T}$ is a $\Phi$-solution for economy $E^{T}$ if for all $t\geq 1$ the following hold:
  1. 1.
    For all $a\in\pazocal{A}_{t}(m^{t-1})$, a matching $\overline{m}\in\varphi_{t}(m^{t-1},a)$ exists such that $U_{t}(a,m)\geq U_{t}(a,\overline{m})$,
  2. 2.
    For all $b\in\pazocal{B}_{t}(m^{t-1})$, a matching $\overline{m}\in\varphi_{t}(m^{t-1},b)$ exists such that $V_{t}(b,m)\geq V_{t}(b,\overline{m})$,
  3. 3.
    No pair $(a,b)\in\pazocal{A}_{t}(m^{t-1})\times\pazocal{B}_{t}(m^{t-1})$ exists such that $u(a,b)>U_{t}(a,m)\text{ and }v(a,b)>V_{t}(b,m)$.
Let $\Sigma_{T}^{\Phi}(E^{T})$ denote the set of $\Phi$-solutions for $E^{T}$.
In words, a matching $m$ is a $\Phi$-solution for $E^{T}$ if in each period the following holds. First, each agent $k$ who is available to match in period $t$ prefers $m$ to their worst conjectured matching (parts 1 and 2 in 6). Note that this implies that if $m_{t}(k)\neq k$, then $k$ prefers their matching partner to waiting to be matched, and if $m_{t}(k)=k$, then $k$ has no objections to the continuation matching, $(m_{s})_{s=t+1}^{T}$.Footnote 5Footnote 5Footnote 5In a static economy, only the first kind of block may exist. The second block only shows up in a dynamic economy (see Doval, 2022 for a discussion). For instance, if the continuation matching is not self-enforcing, an agent may object to it. As we discuss later on, consistency of the conjectures rules is very much related to the second class of blocks. In both cases, we say $k$ prefers to be available to match in period $t$. Second, no pair of agents who can match in period $t$ prefer to match with each other to matching according to $m$ (part 3 in 6). Note that $(a,b)$ being able to match in period $t$ does not imply $a$ and $b$ arrive in the same period, but that under matching $m$ they meet in period $t$ and are able to block $m$.
6 is in the spirit of the static notion of stability:Footnote 6Footnote 6Footnote 66 is also similar to that of dynamic stability in Doval (2022), except that the conjectures in dynamic stability are recursively defined using the solution concept. In other words, 6 nests that of dynamic stability by allowing for arbitrary conjectures. $\Phi$-solutions are defined by the absence of pairwise blocks and the requirement that each agent is matched to a partner that is preferred to remaining unmatched. In contrast to the static notion of stability, the value of remaining unmatched is determined by the conjectures. Indeed, when $T=1$, the correspondence $\Sigma_{1}^{\Phi}$ reduces to the set of stable matchings (4). To see this, consider a length-1 economy, $E^{1}=(A_{1},B_{1})$, a matching $m$ for $E^{1}$, and an agent $k\in A_{1}\cup B_{1}$. Because all matchings $\overline{m}\in\varphi_{1}(m^{0},k)$ are such that $\overline{m}_{1}(k)=k$, then parts 1 and 2 of 6 simply state that $k$ prefers $m$ to remaining single. That is, $m$ is individually rational. Furthermore, part 3 implies that $m$ has no pairwise blocks. Thus, when $T=1$, 6 reduces to 4 for any family of conjectures. We record this observation for future reference:
Observation 1.
For all economies of length $1$, $E^{1}$, and for all conjectures, $\Phi$, the set of $\Phi$-solutions for $E^{1}$ corresponds to the set of stable matchings for $E^{1}$. That is, for all $E^{1}$ and all $\Phi$, $\Sigma_{1}^{\Phi}(E^{1})=\mathrm{S}(E^{1})$.
Among other things, 1 implies that $\Phi$-solutions are always nonempty for economies of length $1$. We use this property repeatedly in our proofs.
Whereas 6 is fairly general, it is worth noting that the definition rules out solution concepts in which the conjectures depend on the matching under consideration beyond the induced economy in period $t$. For instance, when considering whether to block matching $m$ in period $t$, if agent $k$ presumes all other contemporaneous agents–except perhaps for $k$’s matching partner–remain matched according to $m_{t}$, then $k$’s conjectures depend on the outcome specified by $m$ in period $t$. (These conjectures are referred to as passive conjectures; see Chowdhury (2004) and Doval (2015).) Instead, 6 presumes we can define the conjectures without reference to any specific matching. The matching under consideration only matters inasmuch as it determines who can match in each period, and hence whose conjectures one should check in each period.
We close this section by providing a recursive definition of the set of $\Phi$-solutions for a length-$T$ economy, which simplifies the formal statements in the following sections:
Definition 7 (Recursive).
Given the correspondences $(\Sigma_{t}^{\Phi})_{t=1}^{T-1}$, matching $m\in\pazocal{M}_{T}$ is a $\Phi$-solution for economy $E^{T}$ if the following hold:
  1. 1.
    For all $a\in A_{1}$, a matching $\overline{m}\in\varphi_{1}(\emptyset,a)$ exists such that $U_{1}(a,m)\geq U_{1}(a,\overline{m})$,
  2. 2.
    For all $b\in B_{1}$, a matching $\overline{m}\in\varphi_{1}(\emptyset,b)$ exists such that $V_{1}(b,m)\geq V_{1}(b,\overline{m})$,
  3. 3.
    No pair $(a,b)\in A_{1}\cup B_{1}$ exists such that $u(a,b)>U_{1}(a,m)\text{ and }v(a,b)>V_{1}(b,m)$, and
  4. 4.
    $(m_{t})|_{t=2}^{T}\in\Sigma_{T-1}^{\Phi}(E_{2}^{T}(m_{1}))$.Footnote 7Footnote 7Footnote 7We abuse notation slightly. Whereas $\Sigma_{T-1}^{\Phi}(E_{2}^{T}(m_{1}))$ defines a matching only for the agents who can match from $t=2$ onward–i.e., it is an element of $\pazocal{M}_{T-1}(E_{2}^{T}(m_{1}))$$m$ also specifies the outcome for those who have matched in period $1$. Thus, part 4 should be read as “$(m_{t})_{t=2}^{T}$ coincides with an element of $\Sigma_{T-1}^{\Phi}(E^{T}_{2}(m_{1}))$ for the agents who are yet to be matched at the end of period 1, and coincides with $m_{1}$, otherwise.”
7 simply states that matching $m$ is a $\Phi$-solution for $E^{T}$ if the following two conditions hold. First, from period $2$ onwards, $m$ specifies a $\Phi$-solution for the induced economy of length $T-1$, $E_{2}^{T}(m_{1})$. Second, $m$ has no blocks in period $1$ in the sense of 6. Because the correspondence $\Sigma_{T-1}^{\Phi}$ only depends on the conjectures from period $2$ onward, this recursive definition allows us in what follows to focus on the properties of $\varphi_{1}(\emptyset,\cdot)$, taking as given that the set $\Sigma_{T-1}^{\Phi}$ is nonempty.
4 Consistent conjectures
In this section, we present a sufficient condition on the family of conjectures, $\Phi$, such that the set of $\Phi$-solutions is nonempty (Theorem 4.1). We dub the conjectures satisfying the sufficient condition as consistent conjectures. Consistent conjectures guarantee that a candidate matching, defined below, is a solution. In what follows, we first introduce the candidate matching(s), and then we state Theorem 4.1. After discussing the intuition behind the result, we connect it to similar observations in Hafalir (2008) and Doval (2022). We conclude this section by showing that the extension to dynamic matching markets of a standard solution concept in matching with externalities, rational expectations, does not entail consistent conjectures.
Candidate matchings
Given an economy $E^{T}$ and the family of conjectures $\Phi$, we introduce a procedure to construct a set of natural candidates to be $\Phi$-solutions for $E^{T}$. This procedure uses the conjectures to transform the dynamic economy into a sequence of static economies without externalities. The next definition is key:
Definition 8 (One-period economy induced by $\varphi_{1}$).
Given an economy $E^{T}$ and a matching $m\in\pazocal{M}_{T}$, the one-period economy induced by the collection $\{\varphi_{1}(\emptyset,k):k\in A_{1}\cup B_{1}\}$ is the length-1 economy defined as follows:
  1. 1.
    Agents are $A_{1}\cup B_{1}$,
  2. 2.
    Preferences $\{u_{\varphi_{1}}(a,\cdot):B_{1}\cup\{a\}\mapsto\mathbb{R}:a\in A_{1}\},\{v_{\varphi_{1}}(\cdot,b):A_{1}\cup\{b\}\mapsto\mathbb{R}:b\in B_{1}\}$ are given by:
    1. (a)
      For each agent $a\in A_{1}$, $u_{\varphi_{1}}(a,b)\geq u_{\varphi_{1}}(a,a)$ only if $u(a,b)\geq\min\{U_{1}(a,\overline{m}):\overline{m}\in\varphi_{1}(\emptyset,a)\}$. Moreover, for $b,b^{\prime}\in B_{1}$, $u_{\varphi_{1}}(a,b)\geq u_{\varphi_{1}}(a,b^{\prime})$ only if $u(a,b)\geq u(a,b^{\prime})$.
    2. (b)
      For each agent $b\in B_{1}$, $v_{\varphi_{1}}(a,b)\geq v_{\varphi_{1}}(b,b)$ only if $v(a,b)\geq\min\{V_{1}(b,\overline{m}):\overline{m}\in\varphi_{1}(\emptyset,b)\}$. Moreover, for $a,a^{\prime}\in A_{1}$, $v_{\varphi_{1}}(a,b)\geq v_{\varphi_{1}}(a^{\prime},b)$ only if $v(a,b)\geq v(a^{\prime},b)$.
We denote by $E_{1,\varphi_{1}}$ the one-period economy induced by $\varphi_{1}$.
In the one-period economy induced by the conjectures $\varphi_{1}$, agents’ preferences are truncated using their worst conjectured matching. That is, in the one-period economy induced by $\varphi_{1}$, an agent $a$ ranks matching partners on side $B$ according to $u(a,\cdot)$, except that agents on side $B$ who are worse than the worst conjectured continuation matching are now deemed unacceptable. In other words, in the one-period economy induced by $\varphi_{1}$, each agent in $A_{1}\cup B_{1}$ has as an outside option the worst conjectured matching.
The construction in 8 can be understood from the point of view of both matching with externalities and dynamic programming. On the one hand, Sasaki and Toda (1996) show how to use a set of conjectures to construct an economy without externalities (we discuss this construction formally after Theorem 4.1). 8 builds on their insight: We use the conjectures $\varphi_{1}$ to define each period-1 agent’s value from remaining unmatched independently of other contemporaneous agents’ matching outcomes. Thus, no externalities exist in the economy $E_{1,\varphi_{1}}$. On the other hand, 8 is also reminiscent of the logic behind dynamic programming: We can reduce a dynamic economy to a static one by appropriately specifying each agent’s continuation values, which in this case are given by their value of remaining unmatched as defined by $\varphi_{1}$.
Suppose we know that $\Sigma_{T-1}^{\Phi}$ is a nonempty-valued correspondence and consider the following set of candidate matchings:
$M^{\star}(\varphi_{1},\Sigma_{T-1}^{\Phi})=\left\{m^{\star}\in\pazocal{M}_{T}:m^{\star}_{1}\in\mathrm{S}\left(E_{1,\varphi_{1}}\right)\text{and }(m^{\star}_{t})_{t=2}^{T}|_{E_{2}^{T}(m^{\star}_{1})}\in\Sigma_{T-1}^{\Phi}(E_{2}^{T}(m^{\star}_{1}))\right\}.$
In words, a candidate matching, $m^{\star}\in M^{\star}$ satisfies the following. In period $1$, $m_{1}^{\star}$ is a stable matching for $E_{1,\varphi_{1}}$, i.e., $m^{\star}_{1}\in\mathrm{S}\left(E_{1,\varphi_{1}}\right)$. Because the one-period economy induced by $\varphi_{1}$ is devoid of any externalities, its set of stable matchings is nonempty and hence, at least one such matching exists (Gale and Shapley, 1962). Given the period-1 matching, $m^{\star}_{1}$, the continuation matching, $(m^{\star}_{t})_{t=2}^{T}$ is selected from $\Sigma_{T-1}^{\Phi}(E_{2}^{T}(m^{\star}_{1}))$, which is nonempty by assumption. Because $\mathrm{S}\left(E_{1,\varphi_{1}}\right)$ may not be a singleton and given one such stable matching, multiple continuation matchings may exist in the solution set of the induced economy, many candidates matchings may exist.Footnote 8Footnote 8Footnote 8Whereas many stable matchings for the one-period economy induced by the period-1 conjectures may exist, the Lone Wolf theorem (McVitie and Wilson, 1970) implies they all have the same set of unmatched agents and hence, they all induce the same continuation economy.
Any matching in $M^{\star}(\varphi_{1},\Sigma_{T-1}^{\Phi})$ is a natural candidate to be a $\Phi$-solution because it satisfies most of the conditions in 7. To see this, fix any such matching, $m^{\star}$, and note the following. First, $m^{\star}$ satisfies part 4 of 7 by construction. Second, $m^{\star}$ satisfies parts 1, 2, and 3 of 7 for all period-1 agents $k\in A_{1}\cup B_{1}$ such that $m^{\star}_{1}(k)\neq k$. By definition of $E_{1,\varphi_{1}}$, every agent who matches in $t=1$ is matched to someone better than their worst conjectured matching. Furthermore, because $m^{\star}_{1}$ is chosen from $\mathrm{S}\left(E_{1,\varphi_{1}}\right)$, no pair of matched agents under $m^{\star}_{1}$ can form a blocking pair.
Toward consistent conjectures
As we explain next, the only reason any such matching $m^{\star}$ may fail to be a $\Phi$-solution is that a “gap” exists between the conjectures of an agent who is unmatched in $t=1$ under $m^{\star}$ and the matching $m^{\star}$. To see this, note that the above discussion implies that if $m^{\star}$ is not a $\Phi$-solution, then $m^{\star}$ has a block involving a period-$1$ agent, $k$, who is unmatched in $t=1$, i.e., $m^{\star}_{1}(k)=k$. We claim next that if $k$ blocks $m^{\star}$, then $m^{\star}$ is not a conjecture for $k$. To fix ideas, let $k=a\in A_{1}$ and suppose $a$ blocks $m^{\star}_{1}$ with an agent $b\in B_{1}$. Because $m^{\star}_{1}\in\mathrm{S}\left(E_{1,\varphi_{1}}\right)$, it must be that at least one of $a$ or $b$ prefer their worst conjectured matching to matching with each other; assume for simplicity that this holds for $a$. Because $a$ wants to block $m^{\star}$ with $b$, it must be that $a$’s outcome under $m^{\star}$ is worse than the worst matching in $\varphi_{1}(\emptyset,a)$. In other words, when considering what matchings may ensue in the continuation economy, $a$ does not deem $m^{\star}$ a feasible conjecture. Because $a$ only cares about the continuation matching $(m^{\star}_{t})_{t=2}^{T}$, that $m^{\star}$ is not a feasible conjecture is akin to agent $a$’s conjectures not being consistent with the solution concept in the continuation. If instead $a$’s conjectures were consistent with the solution concept in the continuation, then they would not be willing to block $m^{\star}$ with $b$.
The above discussion motivates the result in Theorem 4.1:
Theorem 4.1 (Solutions induced by consistent conjectures are nonempty).
Suppose $\Sigma_{T-1}^{\Phi}$ is a nonempty-valued correspondence. Fix $E^{T}$ and a candidate matching $m^{\star}\in M^{\star}(\varphi_{1},\Sigma_{T-1}^{\Phi})$. Suppose that for all $k\in A_{1}\cup B_{1}$ such that $m^{\star}_{1}(k)=k$, we have that
$m^{\star}\in\varphi_{1}(\emptyset,k).$ (CC)
Then, $m^{\star}\in\Sigma_{T}^{\Phi}(E^{T})$.
The proof of Theorem 4.1 and other results can be found in the appendix.
In words, if for some candidate matching $m^{\star}$, $m^{\star}$ is a conjectured matching for each of the unmatched agents under $m^{\star}$ in period $1$, then $m^{\star}$ is a solution for the economy. Whereas agents who match in period $1$ under $m^{\star}$ may have different conjectures about the matching that would ensue conditional on remaining unmatched, consistency requires that enough commonality exists in the conjectures of those agents who remain unmatched. Indeed, all agents who remain unmatched in period $1$ should agree that $(m^{\star}_{t})_{t=2}^{T}$ is a feasible continuation matching.
As stated, the condition in Theorem 4.1 seems difficult to check without actually computing $m^{\star}$. Yet, Theorem 4.1 suggests a family of conjectures that always lead to nonempty solutions without actually computing $m^{\star}$. Indeed, it follows from Theorem 4.1 that if each period-1 agent conjectures that from $t=2$ onward, the continuation matching is an element of $\Sigma_{T-1}^{\Phi}(\cdot)$, then $\Sigma_{T}^{\Phi}$ is nonempty valued. Formally, define for each $k\in A_{1}\cup B_{1}$,
$\varphi^{*}_{1}(\emptyset,k)=\left\{\overline{m}\in\pazocal{M}_{T}(\emptyset,k):(\overline{m}_{t})_{t=2}^{T}\in\Sigma_{T-1}^{\Phi}(E_{2}^{T}(\overline{m}_{1}))\right\}.$ (4.1)
These conjectures reflect that $k$ believes that the economy abides by the set of $\Phi$-solutions from period $2$ onward. Thus, for each conjectured period-1 matching, $\overline{m}_{1}$, $k$ anticipates that the remaining unmatched agents and the new entrants will object to any matching $(\overline{m}_{t})_{t=2}^{T}$ that does not conform with $\Sigma_{T-1}^{\Phi}(\overline{m}_{1})$.Footnote 9Footnote 9Footnote 9These conjectures also appear in the solution concept for two-sided, repeated matching markets in Kotowski (2019). For instance, if $T=2$, these conjectures rule out matchings $\overline{m}$ such that the period-2 matching, $\overline{m}_{2}$ is not stable in the induced economy, $E_{2}^{2}(\overline{m}_{1})$. Interestingly, that these conjectures induce a nonempty solution concept implies that the perfection requirement alone is compatible with the existence of solutions for dynamic matching markets. Instead, difficulties may arise if the conjectures also place structure on the period-$t$ matching, as we do in Section 5.
Finally, even if the statement of Theorem 4.1 relies on the recursive definition of the set of $\Phi$-solutions for a length-$T$ economy taking the non-emptiness of $\Sigma_{T-1}^{\Phi}$ as given, the procedure to construct $m^{\star}$ and the statement of Theorem 4.1 can be expressed solely in terms of the model primitives. We do this in Appendix A for completeness. Here we note that this only leads to requiring consistency of the conjectures in each period along the path of $m^{\star}$: Given $m^{\star}$, consistency is defined as the requirement that for all $t\geq 1$, all $k\in\pazocal{A}_{t}(m^{\star^{t-1}})\cup\pazocal{B}_{t}(m^{\star^{t-1}})$ such that $m^{\star}_{t}(k)=k$, then $m^{\star}\in\varphi_{t}(m^{\star^{t-1}},k)$ (see Theorem A.1). Similarly, we recursively build the conjectures and the solution concept corresponding to Equation 4.1 and show that a solution concept built off of conjectures that rule out matchings that do not conform with the solution concept in the continuation economy is always nonempty (see A.1 in Appendix A).
We close this section by (i) discussing the connection with Hafalir (2008) and Doval (2022), and why consistent conjectures are conceptually appealing, and (ii) illustrating through an example that rational expectations, a traditional solution concept in matching with externalities, does not have consistent conjectures.
Connection with Hafalir (2008) and Doval (2022)
Hafalir (2008) studies stability in two-sided, one-to-one, matching markets with payoff externalities. The starting point of his analysis is that the negative results in Sasaki and Toda (1996) follow from conjectures being exogenously defined without regard to the economy at hand. The first main result in Hafalir (2008) provides a sufficient condition on the conjectures for a stable matching to exist in his environment. To explain the condition, we review the construction in Hafalir (2008). Using the conjectures, Hafalir (2008) constructs a two-sided economy without externalities as in Sasaki and Toda (1996). In present notation, consider a one-period economy with agents in $A_{1}\cup B_{1}$. For each pair of agents, $(a,b)\in A_{1}\times B_{1}$, a conjecture $\varphi(a,b)$ is a subset of the set of matchings such that $(a,b)$ is matched.Footnote 10Footnote 10Footnote 10Hafalir (2008) does not consider blocks by individual agents. Given the conjectures, say $a$ prefers $b$ over $b^{\prime}$ if $a$ prefers the worst matching in $\varphi$$(a,b)$ over the worst matching in $\varphi(a,b^{\prime})$. The preferences for agents on side $B$ are similarly defined. Under these induced preferences, a stable matching is guaranteed to exist. Fix any such matching, $\overline{m}$. Hafalir (2008, Proposition 2) shows that if for all pairs $(a,b)$ such that $\overline{m}(a)=b$, we have that $\overline{m}\in\varphi(a,b)$, then $\overline{m}$ is stable in the economy with externalities.
Theorem 4.1 extends the logic of the result in Hafalir (2008) to dynamic matching markets. In contrast to Hafalir (2008), no payoff externalities exist in our model; that is, an agent $k$’s payoff from a matching $m$ depends only on the agent’s matching partner, $m_{T}(k)$. Instead, other agents’ outcomes are relevant to define $k$’s payoff from remaining unmatched in a given period. For this reason, Theorem 4.1 only requires that consistency holds for those agents who are unmatched under $m^{\star}$, whereas the condition in Hafalir (2008) must hold for all matched pairs.
Theorem 4.1 identifies a key property that underlies the existence result in Doval (2022). In fact, Remark 2 in that paper points out that the conjectures corresponding to dynamic stability satisfy condition CC for any dynamically stable matching–not just the corresponding candidate matchings $m^{\star}$–and discusses the key role this property plays in the existence result. In Doval (2022), the conjectures are defined recursively based on what the solution concept specifies in the continuation, similar to the conjectures in Equation 4.1. Because our framework isolates the conjectures from the solution concept, it allows us to identify the role of consistency without reference as to how the conjectures relate to the solution concept. In doing so, the result offers an alternative way to construct solution concepts for dynamic matching markets by focusing attention on guaranteeing the conjectures are consistent. Section 5 illustrates this approach.
Why consistent conjectures?
Because Theorem 4.1 offers a recipe of sorts to construct solution concepts in dynamic matching markets with nonempty predictions, considering the conceptual merits of consistent conjectures is natural. We consider two in what follows: The first builds on a similar argument in Hafalir (2008); the second expands on the notion that consistency is akin to the agents understanding and agreeing that the solution concept represents the set of self-enforcing matchings in the continuation economy.
First, it is instructive to consider the implications of not having consistent conjectures in the context of the candidate matchings. To fix ideas, consider one such $m^{\star}$ and suppose an unmatched agent on side $A$, $a$, wishes to block $m^{\star}$ with an agent on side $B$, $b$. Suppose also that $a$ prefers the worst matching in $\varphi_{1}(\emptyset,a)$ to matching with $b$, and prefers matching with agent $b$ to $m^{\star}$, so that $m^{\star}\notin\varphi_{1}(\emptyset,a)$. There is a sense in which $a$’s block is not credible: Whereas $a$ prefers to match with $b$ over $m^{\star}$, as soon as $a$ matches with $b$, $a$ would prefer to break up with $b$. After all, $a$ prefers the worst matching in $\varphi_{1}(\emptyset,a)$ to matching with $b$. Thus, whenever the failure of consistency rules out a candidate matching $m^{\star}$ as a solution, this may be because the solution concept allows for these non-credible blocks.
To make the second point, considering a more general definition of consistency which dispenses with the candidate matchings is useful. Given a family of conjectures, $\Phi$, and the induced solution concept $\{\Sigma_{t}^{\Phi}:t\in\{1,\dots,T\}\}$, we say the solution concept satisfies generalized consistency if the following holds. For all economies $E^{T}$ and $\Phi$-solutions $m$ for $E^{T}$, for all periods $t$, and $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$, if $m_{t}(k)=k$, then $m\in\varphi_{t}(m^{t-1},k)$. That is, we require that condition CC holds not necessarily for the candidate matching $m^{\star}$, which may or may not be a solution, but for the actual solutions (provided they exist).
Suppose we know the set of $\Phi$-solutions is nonempty, but does not satisfy generalized consistency. For concreteness, let $m$ be a solution for $E^{T}$ for which a period $t\geq 1$ and an agent $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$ exist such that $m_{t}(k)\neq k$ and $m\notin\varphi_{t}(m^{t-1},k)$. Because $m$ is a solution, the issue of credible blocks discussed above does not arise. However, that $k$ does not conceive of $m$ as a potential outcome for the continuation economy suggests a tension exists between what the solution concept prescribes in the continuation economy in which $k$ is one of the agents available to match, and what $k$ expects would happen if they would object to $m$. That $m$ is not part of $k$’s conjectures means $k$ believes that $m$ cannot be enforced in the continuation. Thus, if the solution concept captures the set of self-enforcing outcomes in the continuation economy and $k$ agrees with this assessment, then $m$ should be part of $k$’s conjectures and the solution concept should satisfy generalized consistency. As we illustrate below, the rational expectations solution fails both consistency and generalized consistency.
Not all solution concepts have consistent conjectures
It is natural to wonder how permissive consistency is as a requirement. We now illustrate via 4.1 that the conjectures in the rational expectations solution do not satisfy consistency.Footnote 11Footnote 11Footnote 11Because consistency is only a sufficient condition, this is not enough to conclude the rational expectations solution may be empty. In recent work, Nicolò et al. (2024) show the rational expectations solution may be empty in dynamic matching markets. Before introducing the example, it is worth defining the rational expectations solution in the context of the dynamic economy. To that end, fix an economy $E^{T}$ and a period-1 agent, $k\in A_{1}\cup B_{1}$. Define the length-$T$ economy, $E^{T}_{\setminus k}$ to be an economy that coincides with $E^{T}$ except that agent $k$ arrives in period $2$.Footnote 12Footnote 12Footnote 12Formally, if $k\in A_{1}$, $E^{T}_{\setminus k}=(A_{1}\setminus\{k\},B_{1},A_{2}\cup\{k\},B_{2},\dots,A_{T},B_{T})$. Rational expectations presumes we have already defined the rational expectations solutions for $E^{T}_{\setminus k}$, denote this by $\mathrm{RE}_{T}(E^{T}_{\setminus k})$. Under rational expectations, agent $k$ believes that when they decide to be unavailable to match in period $1$, the matching that ensues in the economy from period $1$ onward is governed by the solution concept while respecting $k$’s decision to not participate in the period-1 matching. Formally, agent $k$’s conjectures are given by $\mathrm{RE}_{T}(E^{T}_{\setminus k})$.Footnote 13Footnote 13Footnote 13Note the slight abuse of notation. We defined conjectures in 5 as a subset of $\pazocal{M}_{T}(\emptyset,k)$ so that a conjecture specifies that agent $k$ is unmatched in period $1$. Instead, $\mathrm{RE}_{T}(E^{T}_{\setminus k})$ is a subset of the set of matchings for $E^{T}_{\setminus k}$ in which $k$ only arrives in period $2$ and hence, specifying $k$’s outcome in period $1$ is not necessary.
Example 4.1 (Rational expectations does not satisfy consistency).
Arrivals are given by $A_{1}=\{a_{1},a_{2},a_{3}\},B_{1}=\{b_{1},b_{2}\},A_{2}=\{a_{4}\},B_{2}=\{b_{3},b_{4}\}$. Below, we list the agents’ preferences. If $(b,0)$ appears before $(b^{\prime},1)$ in the ranking of an agent on side $A$, then they prefer to match immediately with $b$ than wait one period to match with $b^{\prime}$. That is, the 00s and $1$s are the exponents of the discount factors, and the list provides the ranking of the discounted utilities $\{\delta_{\cdot}^{t-1}u(\cdot,b):a\in A_{1}\}$. For the $t=1$-agents for whom we present the ranking of utilities alone (i.e., $a_{2},b_{1}$), we assume the discount factor is close to $1$–though not exactly $1$–so that the (static) rankings also represent the rankings over the dynamic matchings.
$\begin{array}[]{lccccc}a_{1}:&(b_{4},0)&(b_{2},0)&(b_{1},0)&(b_{4},1)&\dots\\ a_{2}:&b_{3}&b_{4}&b_{2}&b_{1}&\\ a_{3}:&(b_{3},0)&(b_{2},0)&(b_{3},1)&\dots&\\ a_{4}:&b_{3}&b_{1}&b_{4}&&\end{array}\hskip 28.45274pt\begin{array}[]{lccccc}b_{1}:&a_{4}&a_{1}&a_{2}&&\\ b_{2}:&(a_{2},0)&(a_{2},1)&(a_{3},0)&(a_{1},0)&(a_{3},1)\\ b_{3}:&a_{3}&a_{4}&a_{2}&&\\ b_{4}:&a_{4}&a_{2}&a_{1}&&\end{array}$
In this economy, the following matching, $m^{\star}$, is a candidate matching:
$m^{\star}=\left(\begin{array}[]{lcl}a_{1}&\_\_&b_{1}\\ a_{2}&\_\_&b_{2}\\ \hline\cr a_{3}&\_\_&b_{3}\\ a_{4}&\_\_&b_{4}\end{array}\right).$
Furthermore, $m^{\star}$ is a rational expectations solution for the economy, which is supported by the following conjectures.Footnote 14Footnote 14Footnote 14We provide the supporting details for this example in Appendix A. Agent $a_{2}$ cannot block $m^{\star}$ because
$m_{a_{2}}=\left(\begin{array}[]{lcl}a_{1}&\_\_&b_{1}\\ \hline\cr a_{2}&\_\_&b_{2}\\ a_{3}&\_\_&b_{3}\\ a_{4}&\_\_&b_{4}\end{array}\right),$
is a rational expectations solution in $E^{2}_{\setminus\{a_{2}\}}$. Furthermore, agent $b_{1}$ cannot block because
$m_{b_{1}}=\left(\begin{array}[]{lcl}a_{3}&\_\_&b_{2}\\ \hline\cr a_{1}&\_\_&b_{1}\\ a_{2}&\_\_&b_{4}\\ a_{4}&\_\_&b_{3}\end{array}\right),$
is a rational expectations solution in $E_{\setminus\{b_{1}\}}^{2}$.
However, the conjectures do not satisfy consistency because $m^{\star}$ is not a valid conjecture for $a_{3}$: In every rational expectations solution in $E_{\setminus\{a_{3},b_{1}\}}^{2}$, $b_{1}$ is matched with $a_{4}$, making it impossible to support a solution in $E_{\setminus\{a_{3}\}}^{2}$ which has $b_{1}$ matched with $a_{1}$. Because $m^{\star}$ is a rational expectations solution for the economy, it follows that the solution concept also fails generalized consistency.
The failure of consistency in the rational expectations solution can be tied back to the timing assumptions implicit in the solution concept and the conjectures. Given their conjectures, the solution concept presumes the $t=1$-agents decide simultaneously whether to be available to match in $t=1$. Instead, each agent’s conjectures presume they are the first to decide not to be available to match in $t=1$, while the remaining contemporaneous agents remain available to match in $t=1$. For instance, when $b_{1}$ considers not being available to match in $t=1$, $b_{1}$ presumes $a_{3}$ is available to match in $t=1$–and in fact, $a_{3}$ must match in $t=1$ to prevent $b_{1}$ from blocking $m^{\star}$. Instead, $a_{3}$’s conjectures presume $a_{3}$ cannot match in $t=1$, and given this, $b_{1}$ is able to match with $a_{4}$ in $t=2$. For this reason, $m^{\star}$ cannot be compatible with $a_{3}$’s conjectures: supporting $m^{\star}$ as a rational expectations solution requires that $a_{3}$ may possibly match in $t=1$.
The analysis so far identified a property of the family of conjectures, $\Phi$, that guarantees existence, implicitly separating what the solution concept describes “on path” from the conjectures that support this path. Instead, we have said little about what these conjectures entail. For instance, whereas consistency implies the conjectures are consistent with the solution concept “on the path” (i.e., for the agents who remain unmatched), it does not require that the same holds “off the path” (i.e., for the agents who are matched). Consequently, agents could be ascribing to $m^{\star}$ for fear of facing continuation matchings that are, in some sense, not credible; a property that would make the $\Phi$-solutions less appealing. Conjectures such as those in Equation 4.1 satisfy consistency and avoid these “non-credible” threats. However, they impose no structure on the matching among the period-1 agents. Section 5 introduce two families of conjectures (and their corresponding solution concepts) addressing these issues.
5 Two solution concepts with consistent conjectures
In this section, we introduce two families of consistent conjectures and their corresponding solution concepts. Section 5.1 defines CVR-dynamic stability, a refinement of dynamic stability in Doval (2022). Section 5.2 defines sophisticated dynamic stability, an extension to dynamic matching markets of the eponymous solution concept in Hafalir (2008).
5.1 Continuation-values-respecting dynamic stability
The solution concept we introduce in this section, CVR-dynamic stability, builds on the conjectures in Equation 4.1. Whereas the conjectures in Equation 4.1 require agents to believe that the continuation matching abides by the solution concept, they place no restriction on what period-$t$ matching an agent expects would ensue conditional on not matching in period $t$. CVR-dynamic stability imposes a minimal requirement on the period-$t$ matching that reflects other agents’ continuation values. In what follows, we first define the requirement on the period-$t$ matching (9), the conjectures for CVR-dynamic stability (see Equation 5.3), and the solution concept (10). We then show in 5.1 that CVR-dynamic stability has consistent conjectures. We end the section by discussing the connection with dynamic stability in Doval (2022).
To define CVR-dynamic stability, we first introduce a condition on the period-$t$ matching. To that end, fix an economy, $E^{T}$, and a matching through period $t-1$, $m^{t-1}$. Suppose that for each agent $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$, we are given a set of matchings $\varphi_{k}\subset\pazocal{M}_{T}(m^{t-1},k)$. Similar to 8, we can define the one-period economy with agents in $\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$ induced by $\{\varphi_{k}:k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})\}$, which we denote by $E_{t,\varphi}$. Given a period-$t$ matching, $m_{t}$, let $E_{t,\varphi}^{m_{t}}$ denote the restriction of $E_{t,\varphi}$ to the set of agents $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$ such that $m_{t}(k)\neq k$.
Definition 9 (Stability among those who match relative to continuation values).
The period-$t$ matching, $m_{t}$, is stable among those who match in period $t$ relative to $\varphi$ if $m_{t}\in\mathrm{S}(E_{t,\varphi}^{m_{t}})$.
As the name suggests, 9 requires the period-$t$ matching satisfies a condition analogous to the static notion of stability among the agents who exit the economy in period $t$. We say analogous because rather than requiring that the period-$t$ matching be individually rational, 9 requires each agent $k$ be matched to a partner they prefer to the worst matching in $\varphi_{k}$.Footnote 15Footnote 15Footnote 15 If for each agent $k$, we take $\varphi_{k}$ to be the subset of individually rational matchings for $k$ in $\pazocal{M}_{T}(m^{t-1},k)$, we obtain the condition in Doval (2022, Definition 4). We discuss this in detail below.
Given 9, we now define the conjectures associated with CVR-dynamic stability. Suppose we have already defined CVR-dynamic stability for economies of length $\{1,\dots,T-t\}$ and let $\hat{\mathrm{DS}}_{T-t}$ denote the correspondence of $\mathrm{CVR}$-dynamically stable matchings for length $T-t$ economies. Fix agent $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$ and define their conjectures under CVR-dynamic stability as follows:
$\varphi_{\hat{\mathrm{DS}},t}(m^{t-1},k)=\left\{\overline{m}\in\pazocal{M}_{T}(m^{t-1},k):\begin{array}[]{ll}\text{ (i) }\overline{m}_{t}\in\mathrm{S}\left(E_{t,\varphi_{\hat{\mathrm{DS}},t}}^{\overline{m}_{t}}\right)\\ \text{ (ii) }(\overline{m}_{s})_{s=t+1}^{T}\in\hat{\mathrm{DS}}_{T-t}(E^{T}_{t+1}(\overline{m}^{t}))\end{array}\right\}.$ (5.3)
That is, when considering not being available to match in period $t$, $k$’s conjectured matchings satisfy two properties. First, like the conjectures in Equation 4.1, $k$ assumes that the continuation matching will be a $\mathrm{CVR}$-dynamically stable matching. Second, $k$ conjectures that those agents who exit the economy in period $t$ under $\overline{m}_{t}$ could not have found a better matching among themselves. That is, $\overline{m}_{t}$ cannot be improved by pairwise blocks, and no agent who matches in period $t$ matches with someone worse than their implicit continuation value according to their own conjectures, $\varphi_{\hat{\mathrm{DS}},t}$.Footnote 16Footnote 16Footnote 16The last part is consistent with the idea that agents simultaneously decide whether to match or remain unmatched given their set of conjectures (recall the discussion in 4.1). This last aspect of $k$’s conjectures is the sense in which they respect other contemporaneous agents’ continuation values (as defined by their conjectures). Despite the fixed-point flavor of the set in Equation 5.3, agent $k$’s conjectures are well-defined: Because $k$ is unmatched in period $t$ under $\overline{m}$, the set $\mathrm{S}\left(E_{t,\varphi_{\hat{\mathrm{DS}},t}}^{\overline{m}_{t}}\right)$ can be defined independently of agent $k$’s conjectures.
We can define the set of $\mathrm{CVR}$-dynamically stable matchings for $E^{T}_{t}(m^{t-1})$, $\hat{\mathrm{DS}}_{T-(t-1)}(E^{T}_{t}(m^{t-1}))$, to be the set of matchings for this economy that satisfy 7 with $\varphi_{\hat{\mathrm{DS}},t}(m^{t-1},\cdot)$ as the conjectures and $\hat{\mathrm{DS}}_{T-t}$ as the continuation solution concept. Working recursively in this way, we can define the family of conjectures $\Phi_{\hat{\mathrm{DS}}}$ for $E^{T}$ to be the collection $\{\varphi_{\hat{\mathrm{DS}},t}(m^{t-1},k):t\geq 1,m^{t-1},k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})\}$. The definition of CVR-dynamic stability is now immediate:
Definition 10 (CVR-dynamic stability).
A matching $m$ for $E^{T}$ is $\mathrm{CVR}$-dynamically stable if it is a $\Phi_{\hat{\mathrm{DS}}}$-solution for $E^{T}$. We denote by $\hat{\mathrm{DS}}_{T}(E^{T})$ the set of such solutions.
5.1 is the main result of this section:
Proposition 5.1.
The family $\Phi_{\hat{\mathrm{DS}}}$ satisfies consistency. Consequently, for all $T\in\mathbb{N}$, the correspondence $\hat{\mathrm{DS}}_{T}$ is nonempty valued.
The proof of 5.1 is in Appendix B. The proof proceeds by induction on the length of the economy, first showing that the set of conjectures is nonempty and then showing they satisfy consistency; Theorem 4.1 then implies the set of $\mathrm{CVR}$-dynamically stable matchings is nonempty. Thus, the key step is showing that the conjectures are nonempty: Because of the fixed-point character of the conjectures in period $t$, we build a recursion whose iteration delivers the set $\varphi_{\hat{\mathrm{DS}},t}$ as a limit.
Comparison with dynamic stability (Doval, 2022)
CVR-dynamic stability is a refinement of dynamic stability, the stability notion in Doval (2022). In particular, if we denote by $\varphi_{\mathrm{DS}}$ the conjectures corresponding to dynamic stability, one can show that for all $t$, all $m^{t-1}$, and $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$, the conjectures $\varphi_{\hat{\mathrm{DS}},t}(m^{t-1},k)\subseteq\varphi_{\mathrm{DS},t}(m^{t-1},k)$, which in turn implies the set of $\mathrm{CVR}$-dynamically stable matchings is a subset of those that are dynamically stable.
To understand the difference between the solution concepts, it is instructive to compare the conjectures in Equation 5.3 to those in Doval (2022). Whereas both solution concepts require agents to conjecture that the continuation matching conforms with the solution concept in the induced continuation economy (i.e., requirement (ii) in Equation 5.3), the conjectures in Doval (2022) impose a weaker requirement on the period-$t$ matching. Namely, that among the agents who exit the economy in period $t$, the period-$t$ matching is individually rational and no blocking pairs exist.Footnote 17Footnote 17Footnote 17Formally, let $\varphi_{k}^{IR}$ denote the set of matchings in $\pazocal{M}_{T}(m^{t-1},k)$ that are individually rational for $k$. Then, in dynamic stability, matching $\overline{m}\in\pazocal{M}_{T}(m^{t-1},k)$ is a conjecture for $k$ if (i) $\overline{m}_{t}$ is stable among those who match relative to continuation values $\varphi_{\cdot}^{IR}$ and (ii) the continuation matching is dynamically stable. In other words, under dynamic stability, an agent $k$ conjectures that other contemporaneous agents are willing to exit in period $t$ with partners that dominate being unmatched through period $T$, even if they may not dominate the worst conjectured matching under $\varphi_{\mathrm{DS},t}(m^{t-1},\cdot)$. At the same time, when considering whether to block the proposed matching in period $t$, the agents available to match in period $t$ compare their payoff under the proposed matching to their worst conjectured matching under $\varphi_{\mathrm{DS},t}(m^{t-1},\cdot)$. That is, in dynamic stability, agents make their decisions to be available to match in period $t$ relative to their conjectures, but these conjectures do not take into account that other contemporaneous agents use their own conjectures when evaluating whether to match in period $t$. CVR-dynamic stability eliminates this gap by requiring the period-$t$ agents to internalize that other contemporaneous agents’ continuation values are exactly those prescribed by their conjectures.
We conclude this section illustrating that CVR-dynamic stability is a strict refinement of dynamic stability:
Example 5.1 ($\hat{\mathrm{DS}}\subsetneq\mathrm{DS}$).
This example is due to Antonio Nicolò, Pietro Salmaso, and Riccardo Saulle. Arrivals are given by $A_{1}=\{a_{1},a_{2}\},B_{1}=\{b_{1},b_{2}\},A_{2}=\{a_{3},a_{4}\},B_{2}=\{b_{3},b_{4}\}$. Assume the discount factors of the period-1 agents are close to $1$, so that the (static) rankings below are enough to describe the agents’ preferences over dynamic matchings.
$\begin{array}[]{lllll}a_{1}:&b_{3}&b_{4}&b_{1}&\\ a_{2}:&b_{4}&b_{2}&b_{3}&\\ a_{3}:&b_{1}&b_{4}&b_{2}&\\ a_{4}:&b_{1}&b_{4}&b_{2}&b_{3}\end{array}\qquad\qquad\qquad\begin{array}[]{lllll}b_{1}:&a_{3}&a_{4}&a_{1}&\\ b_{2}:&a_{2}&a_{3}&a_{4}&\\ b_{3}:&a_{1}&a_{2}&a_{4}&\\ b_{4}:&a_{1}&a_{3}&a_{2}&a_{4}\end{array}$
Consider the following three matchings:
$m^{L}=\left(\begin{array}[]{lcl}a_{2}&\_&b_{2}\\ \hline\cr a_{1}&\_&b_{3}\\ a_{3}&\_&b_{1}\\ a_{4}&\_&b_{4}\end{array}\right)\quad\quad m^{C}=\left(\begin{array}[]{lcl}a_{1}&\_&b_{1}\\ \hline\cr a_{2}&\_&b_{2}\\ a_{3}&\_&b_{4}\\ a_{4}&\_&b_{3}\end{array}\right)\quad\quad m^{R}=\left(\begin{array}[]{lcl}&\emptyset&\\ \hline\cr a_{1}&\_&b_{3}\\ a_{2}&\_&b_{4}\\ a_{3}&\_&b_{1}\\ a_{4}&\_&b_{2}\end{array}\right).$
The matching on the left, $m^{L}$, is a dynamically stable matching, with $m^{C}$ describing the conjectured matching that dissuades $a_{2}$ from waiting to be matched in $t=2$. Importantly, under $m^{C}$, $a_{1}$ and $b_{1}$ match in period $1$. However, it is easy to see that $a_{1}$ can always guarantee matching with $b_{3}$ (and similarly, $b_{1}$ with $a_{3}$) by remaining unmatched in $t=1$. In other words, $m^{C}$ cannot be a conjecture for $a_{2}$ under $\varphi_{\hat{\mathrm{DS}}}$ and hence, $m^{L}$ is not $\mathrm{CVR}$-dynamically stable. Instead, $m^{R}$ is $\mathrm{CVR}$-dynamically stable (and hence, dynamically stable).
5.2 Sophisticated dynamic stability
In this section, we provide the natural extension of the solution concept in Hafalir (2008) to the dynamic environment we study. Because Hafalir (2008) names the solution concept sophisticated expectations, we denote ours by sophisticated dynamic stability. Like CVR-dynamic stability, sophisticated dynamic stability is defined recursively. In what follows, we denote by $\mathrm{SDS}_{t}$ the correspondence that maps economies of length $t$ to their set of sophisticated dynamically stable matchings.
Sophisticated dynamic stability builds on two elements: The first is the rational expectations solution discussed at the end of Section 4. Recall that when agent $k$ chooses to remain unmatched in period $t$, it is as if $k$ induces a new economy from period $t$ onward, where everything is as before, except that now $k$ arrives in period $t+1$. Like the rational expectations solution, sophisticated dynamic stability requires that $k$ never rules out matchings $\overline{m}$ such that from period $t$ onward $\overline{m}$ induces a sophisticated dynamically stable matching in the economy induced by $k$ remaining unmatched in period $t$ (see Equation 5.4 below). The second element is a procedure that guarantees the consistency of the conjectures in sophisticated dynamic stability. Recall that 4.1 shows that the conjectures in the rational expectations solution may fail consistency. The procedure below recursively expands the set of conjectures to guarantee their consistency.
We now proceed to define the conjectures for sophisticated dynamic stability, which we denote by $\varphi_{\mathrm{SDS},t}$. Fix an economy $E^{T}$, and a matching through period $t-1$, $m^{t-1}$. Recall that this induces a continuation economy of length $T-(t-1)$, $E^{T}_{t}(m^{t-1})$. Suppose that for all $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$, we have defined the set of sophisticated dynamically stable matchings for the economy of length $T-(t-1)$ induced by $k$ arriving in period $t+1$, $\mathrm{SDS}_{T-(t-1)}\left(E_{t}^{T}(m^{t-1})_{\setminus k}\right)$.
For each $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$, the conjectures in sophisticated dynamic stability are defined recursively, starting from the following set:
$F_{0}(E^{T}_{t}(m^{t-1}),k)=\left\{\overline{m}\in\pazocal{M}_{T}(m^{t-1},k):(\overline{m}_{s})_{s=t}^{T}\in\mathrm{SDS}_{T-(t-1)}(E_{t}^{T}(m^{t-1})_{\setminus k})\right\}.$ (5.4)
In words, the set $F_{0}(E^{T}_{t}(m^{t-1}),k)$ consists of all sophisticated dynamically stable matchings for the $T-(t-1)$ economy in which $k$ arrives in period $t+1$. We make two observations. First, $F_{0}(E^{T}_{t}(m^{t-1}),k)$ corresponds to the set of conjectures in the “rational expectations solution” using $\mathrm{SDS}_{T-(t-1)}$. Second, in contrast to the notation so far, Equation 5.4 makes the dependence of the conjectures on the continuation economy explicit. The reason is that the set of sophisticated dynamically stable matchings for $E^{T}_{t}(m^{t-1})$ depends on the set of sophisticated dynamically stable matchings for $E^{T}_{t}(m^{t-1})_{\setminus k}$, which in turn depends on the conjectures $\varphi_{\mathrm{SDS},t}(E^{T}_{t}(m^{t-1})_{\setminus k},k^{\prime})$ for $k^{\prime}\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})\setminus\{k\}$, and so on. Thus, keeping track of the economy from period $t$ onwards is helpful for clarity.
Given $F_{0}(E^{T}_{t}(m^{t-1}),\cdot)$, let $\mathrm{S}(E_{t,F_{0}})$ denote the set of stable matchings in the one-period economy induced by $\{F_{0}(E^{T}_{t}(m^{t-1}),\cdot):k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})\}$. Consider now the following set:
$M^{\star}\left(m^{t-1},F_{0},\mathrm{SDS}_{T-t}\right)=\left\{\overline{m}\in\pazocal{M}_{T}(m^{t-1}):\begin{array}[]{l}\text{(i) }\overline{m}_{t}\in\mathrm{S}(E_{t,F_{0}})\\ \text{(ii) }(\overline{m}_{s})_{s=t+1}^{T}\in\mathrm{SDS}_{T-t}(E_{t+1}^{T}(\overline{m}^{t}))\end{array}\right\}.$
$M^{\star}\left(m^{t-1},F_{0},\mathrm{SDS}_{T-t}\right)$ corresponds to the set of matchings such that (i) in period $t$, agents form a stable matching in the one-shot economy induced by their continuation values $F_{0}(E^{T}_{t}(m^{t-1}),\cdot)$, and (ii) the matching from period $t+1$ onward is sophisticated dynamically stable. In other words, $M^{\star}\left(m^{t-1},F_{0},\mathrm{SDS}_{T-t}\right)$ is the set of candidate matchings from Section 4, when the current conjectures are $F_{0}$ and the continuation solution concept is given by $\mathrm{SDS}_{T-t}$.
Recursively, define for $n\geq 1$ and $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$,
$F_{n}(E^{T}_{t}(m^{t-1}),k)=F_{n-1}(E^{T}_{t}(m^{t-1}),k)\cup\left\{\overline{m}\in M^{\star}(m^{t-1},F_{n-1},\mathrm{SDS}_{T-1}):\overline{m}_{t}(k)=k\right\}.$ (5.5)
This process expands the set of conjectures of those agents who find themselves unmatched in the economy with continuation values in $F_{n-1}(E^{T}_{t}(m^{t-1}),\cdot)$. Note that matchings are only added if the elements in $M^{\star}(m^{t-1},F_{n-1},\mathrm{SDS}_{T-1})$ are not part of the set of conjectures $F_{n-1}$.
The recursion in Equation 5.5 is meant to solve the issue discussed in Section 4: If a candidate matching $m^{\star}$ given current conjectures $F_{n-1}$ and the continuation solution concept $\mathrm{SDS}_{T-t}$ is not a solution, then it is blocked by a pair of agents, where at least one of them prefer their worst conjectured matching to matching together. By adding the matchings in $M^{\star}(m^{t-1},F_{n-1},\mathrm{SDS}_{T-1})$, we ensure agents do not dismiss matchings like $m^{\star}$, whose blocks are not credible.
We define $k$’s conjectures as the limit of the set $F_{n}(E^{T}_{t}(m^{t-1}),k)$, that is:
$\varphi_{\mathrm{SDS},t}(E^{T}_{t}(m^{t-1}),k)=\lim_{n\rightarrow\infty}F_{n}(E^{T}_{t}(m^{t-1}),k).$ (5.6)
The set in Equation 5.6 is well defined because the sequence $F_{n}$ is increasing (in the set-inclusion sense) and bounded above by $\pazocal{M}_{T}(m^{t-1},k)$.
We can define the set of sophisticated dynamically stable matchings for $E^{T}_{t}(m^{t-1})$,
$\mathrm{SDS}_{T-(t-1)}(E^{T}_{t}(m^{t-1}))$, to be the set of matchings for this economy that satisfy 7 with $\varphi_{\mathrm{SDS},t}(E^{T}_{t}(m^{t-1}),\cdot)$ as the conjectures and $\mathrm{SDS}_{T-t}$ as the continuation solution concept.Footnote 18Footnote 18Footnote 18To be precise, for a given subset $C\subseteq\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$, we should define the set of sophisticated dynamically stable matchings for $E^{T}_{t}(m^{t-1})_{\setminus C}$ having defined $\mathrm{SDS}_{T-(t-1)}(E^{T}_{t}(m^{t-1})_{\setminus C^{\prime}})$ for all subsets $C^{\prime}\subseteq(\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1}))\setminus C$, where $E^{T}_{t}(m^{t-1})_{\setminus C}$ is the length-$T-(t-1)$ economy in which the agents in $C$ arrive in period $t+1$. However, $E^{T}_{t}(m^{t-1}){\setminus C}$ is yet another economy of length $T-(t-1)$, so this omission is hopefully innocuous. Working recursively in this way, we can define the family of conjectures $\Phi_{\mathrm{SDS}}$ for $E^{T}$. The definition of sophisticated dynamic stability is now immediate:
Definition 11.
A matching $m$ for $E^{T}$ is sophisticated dynamically stable if it is a $\Phi_{\mathrm{SDS}}$-solution for $E^{T}$. We denote by $\mathrm{SDS}_{T}(E^{T})$ the set of such solutions.
Because the conjectures in sophisticated dynamic stability are constructed so that they satisfy consistency, the following result is an immediate corollary of Theorem 4.1:
Corollary 5.1.
For all $T\in\mathbb{N}$, $\mathrm{SDS}_{T}$ is a nonempty-valued correspondence.
6 Conclusions
In this paper, we provide a unifying framework to study existence properties of solution concepts for two-sided, one-to-one, dynamic matching markets in which matching is irreversible. Following the literature on matching with externalities (Sasaki and Toda, 1996; Hafalir, 2008), we give centerstage to agents’ conjectures about the matching that would ensue should they remain unmatched. Given a collection of conjectures, we define a solution concept given these conjectures. We identify a sufficient condition on the conjectures for the corresponding solution concept to be nonempty (Theorem 4.1). Armed with this result, we propose two families of conjectures and, consequently, two new solution concepts for dynamic matching markets with nonempty solutions.
We see several avenues worth exploring and left for future work. First, whereas the rational expectations solution does not satisfy consistency, Nicolò et al. (2023) show that existence is guaranteed in dynamic, one-to-one, one-sided matching markets. Because the same externality is present in both two-sided and one-sided dynamic matching markets, understanding what sets apart these two environments is useful.
Second, whereas we propose two new solution concepts for dynamic matching markets, these are by no means the only solution concepts that satisfy consistency. Further understanding the set of solution concepts that satisfy consistency would allow us to delineate the limits of the predictions afforded by such solution concepts. Nicolò et al. (2024) is a step in this direction.
Finally, because important dynamic matching applications, like sequential assignment in school choice, involve many-to-one matching, extending our results to many-to-one matching markets is natural. As the work of Altinok (2019) and Liu (2023) highlights, new issues arise in the study of such markets and understanding what the analogue of consistency is in these markets is worth exploring.
Appendix A Omitted statements and proofs from Section 4
For completeness, we collect in this section the definitions analogous to those in Section 4 that allow us to describe the construction of the candidate matchings and state Theorem 4.1 solely in terms of the model primitives.
Non-recursive construction of the candidate matchings
As in the main text, consider the family of conjectures $\Phi$.
Definition 12 (One-period economy induced by $\varphi_{t}$ at $m^{t-1}$).
Given an economy $E^{T}$, a period $t$, and a matching through period $t-1$, $m^{t-1}$. The one-period economy induced by the collection $\{\varphi_{t}(m^{t-1},k):k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})\}$ is the length-1 economy defined as follows:
  1. 1.
    Agents are $\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$,
  2. 2.
    Preferences $\{u_{\varphi_{t}}(a,\cdot):\pazocal{B}_{t}(m^{t-1})\cup\{a\}\mapsto\mathbb{R}:a\in\pazocal{A}_{t}(m^{t-1})\},\{v_{\varphi_{t}}(\cdot,b):\pazocal{A}_{t}(m^{t-1})\cup\{b\}\mapsto\mathbb{R}:b\in\pazocal{B}_{t}(m^{t-1})\}$ are given by:
    1. (a)
      For each agent $a\in\pazocal{A}_{t}(m^{t-1})$, $u_{\varphi_{t}}(a,b)\geq u_{\varphi_{t}}(a,a)$ only if $u(a,b)\geq\min\{U_{t}(a,\overline{m}):\overline{m}\in\varphi_{t}(m^{t-1},a)\}$. Moreover, for $b,b^{\prime}\in\pazocal{B}_{t}(m^{t-1})$, $u_{\varphi_{t}}(a,b)\geq u_{\varphi_{t}}(a,b^{\prime})$ only if $u(a,b)\geq u(a,b^{\prime})$.
    2. (b)
      For each agent $b\in\pazocal{B}_{t}(m^{t-1})$, $v_{\varphi_{t}}(a,b)\geq v_{\varphi_{t}}(b,b)$ only if $v(a,b)\geq\min\{V_{t}(b,\overline{m}):\overline{m}\in\varphi_{t}(m^{t-1},b)\}$. Moreover, for $a,a^{\prime}\in\pazocal{A}_{t}(m^{t-1})$, $v_{\varphi_{t}}(a,b)\geq v_{\varphi_{t}}(a^{\prime},b)$ only if $v(a,b)\geq v(a^{\prime},b)$.
We denote by $E_{t,\varphi_{t}}(m^{t-1})$ the one-period economy induced by $\varphi_{t}$ at $m^{t-1}$.
We can now provide the non-recursive definition of the candidate matchings for $E^{T}$ given $\Phi$:
$M^{\star}(\Phi)=\left\{m^{\star}\in\pazocal{M}_{T}:(\forall t\in\{1,\dots,T\})m^{\star}_{t}|_{E_{t,\varphi_{t}}(m^{\star^{t-1}})}\in\mathrm{S}\left(E_{t,\varphi_{t}}(m^{\star^{t-1}})\right)\right\},$
where in a slight abuse of notation we use the notation $M^{\star}(\Phi)$ to signify the candidate matchings that can be constructed solely as a function of the family of conjectures, $\Phi$, for economy $E^{T}$.
The analogous statement to that in Theorem 4.1 now follows:
Theorem A.1.
Fix $E^{T}$ and a matching $m^{\star}\in M^{\star}(\Phi)$. Suppose that for all $t\geq 1$ and all $k$ such that $m^{\star}_{t}(k)=k$, we have that
$m^{\star}\in\varphi_{t}(m^{\star^{t-1}},k).$ (A.1)
Then, for all $t\geq 1$, $(m_{s}^{\star})_{s=t}^{T}|_{E^{T}_{t}(m^{\star^{t-1}})}\in\Sigma_{T-t+1}^{\Phi}(E^{T}_{t}(m^{\star^{t-1}}))$.
Even though the argument leading to the statement of Theorem 4.1 basically provides the proof of that theorem and also of Theorem A.1, we provide a self-contained proof in the notation in this appendix for completeness:
Proof of Theorem A.1.
Assume that (A.1) holds. Note that for $t=T$ the result is vacuously true. Suppose we have established the result for $s\in\{t+1,\dots,T\}$. We show it holds for $s=t$. Note that $m^{\star}_{t}$ immediately satisfies parts 1 and 2 of 6 whenever $m^{\star}_{t}(k)\neq k$. Similarly, it satisfies part 3 of 6 for pairs $(a,b)\in\pazocal{A}_{t}(m^{t-1})\times\pazocal{B}_{t}(m^{t-1})$ such that $m^{\star}_{t}(a)\neq a,m^{\star}_{t}(b)\neq b$.
Consider now $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$ such that $m^{\star}_{t}(k)=k$ and note that (A.1) implies that under $m^{\star}$, $k$ is getting at least the payoff of $k$’s worst conjectured matching under $\varphi_{t}(m^{\star^{t-1}},\cdot)$.
Finally, suppose $m^{\star}_{t}$ fails part 3 of 6 for some pair $(a,b)$ in $\pazocal{A}_{t}(m^{\star^{t-1}})\times\pazocal{B}_{t}(m^{\star^{t-1}})$. Then at least one of the members of the pair must be single under $m^{\star}_{t}$. Without loss of generality, assume $m^{\star}_{t}(a)=a$. Then, $u(a,b)>U_{t}(a,m^{\star})\geq\min\{U_{t}(a,\overline{m}):\overline{m}\in\varphi_{t}(m^{\star^{t-1}},a)\}$, where the second inequality follows from $m^{\star}\in\varphi_{t}(m^{\star^{t-1}},a)$. This chain of inequalities contradicts the definition of $m^{\star}_{t}$. Thus $(m_{s}^{\star})_{s=t}^{T}\in\Sigma_{T-t+1}^{\Phi}(E^{T}_{t}(m^{\star^{t-1}}))$ and the result follows. ∎
Solution concept based on the conjectures in Equation 4.1
In Section 4, we argued that the period-1 conjectures defined in Equation 4.1 lead to nonempty solutions for $E^{T}$ provided that the solution concept for the continuation economy, $\Sigma_{T-1}^{\Phi}$ is also nonempty. For completeness, we now build recursively the conjectures and the solution concept to complete the argument: A solution concept as in 6 built off of conjectures that rule out matchings that are not consistent with the solution concept in the continuation is always nonempty. In what follows, we keep the notation $\varphi^{*}$ for the conjectures and denote by $\Sigma^{\Phi^{*}}$ the induced solution concept.
Fix an economy of length $T$. Recall that 1 implies that no matter what the conjectures are, the set of $\Phi$-solutions for an economy of length $1$ coincide with the set of stable matchings. Given this, fix a matching through period $T-2$, $m^{T-2}$. For $k\in\pazocal{A}_{T-1}\left(m^{T-2}\right)\cup\pazocal{B}_{T-1}\left(m^{T-2}\right)$, define $k$’s conjectures as follows:
$\varphi^{*}_{T-1}(m^{T-2},k)=\left\{\overline{m}\in\pazocal{M}_{T}(m^{T-2},k):\overline{m}_{T}\in\mathrm{S}\left(E^{T}_{T}(\overline{m}^{T-1})\right)\right\}.$
Let $\Sigma^{\Phi^{*}}_{T-1}\left(E^{T}_{T-1}\left(m^{T-2}\right)\right)$ denote the set of $\varphi^{*}_{T-1}$-solutions for $E^{T}_{T-1}\left(m^{T-2}\right)$.
Recursively, suppose we have defined $\Sigma^{\Phi^{*}}_{T-t}$ for economies of length $T-t$, and fix a matching through period $t$, $m^{t-1}$. For $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$, $k$’s conjectures are given by:
$\varphi^{*}_{t}(m^{t-1},k)=\left\{\overline{m}\in\pazocal{M}_{T}(m^{t-1},k):(\overline{m}_{s})_{s=t+1}^{T}\in\Sigma^{\Phi^{*}}_{T-t}\left(E^{T}_{t+1}(\overline{m}^{t})\right)\right\},$
and let $\Sigma^{\Phi^{*}}_{T-t+1}\left(E^{T}_{t}(m^{t-1})\right)$ denote the set of solutions given the family $\Phi^{*}_{\geq t}=\{\varphi^{*}_{s}((m^{t-1},\overline{m}_{t}^{s-1}),k):s\geq t,\overline{m}_{t}^{s-1},k\in\pazocal{A}_{s}(m^{t-1},m_{t}^{s-1})\cup\pazocal{B}_{s}(m^{t-1},m_{t}^{s-1})\}$.
The following statement is an immediate corollary of Theorem 4.1 and the definition of the conjectures $\varphi^{*}$:
Corollary A.1.
For all $T\in\mathbb{N}$, $\Sigma^{\Phi^{*}}_{T}$ is a nonempty-valued correspondence.
A.1 then shows that existence issues for solution concepts in dynamic matching markets do not arise solely from the perfection requirement–i.e., that the continuation matching also abides by the solution concept–but from restrictions the conjectures place on what period-$t$ matchings are feasible.
Omitted details from 4.1
In what follows, we provide supporting details for the derivation in 4.1. First, note that should no one match in period $1$, a unique stable matching exists in $t=2$, which leads to the matching:
$m_{\emptyset}=\left(\begin{array}[]{lcl}&\emptyset&\\ \hline\cr a_{1}&\_\_&b_{2}\\ a_{2}&\_\_&b_{4}\\ a_{3}&\_\_&b_{3}\\ a_{4}&\_\_&b_{1}\end{array}\right).$
It follows that $m_{\emptyset}$ is the unique rational expectations solution in any two-period economy in which at least $b_{1},b_{2}$ are part of the $t=2$-arrivals.
Second, consider $E_{\setminus\{b_{1}\}}^{2}$. We argue that $m_{b_{1}}$ is a rational expectations solution. To see this, note that $b_{2}$ does not wish to block with $a_{1}$, and $a_{3}$ cannot do better than by matching in $t=1$ with $b_{2}$. The only possibility is that $b_{2}$ waits to be matched. However, $m_{\emptyset}$ is a valid conjecture for $b_{2}$ in $E_{\setminus\{b_{1},b_{2}\}}^{2}$, and $b_{2}$ prefers $a_{3}$ to matching with $a_{1}$ in $t=2$.
Third, consider $E_{\setminus\{a_{3},b_{1}\}}^{2}$. We show that $b_{1}$ is always matched to $a_{4}$ in any rational expectations solution and hence $b_{1}$ cannot match with $a_{1}$ in a rational expectations solution for $E_{\setminus\{a_{3}\}}^{2}$. In $E_{\setminus\{a_{3},b_{1}\}}^{2}$, three possibilities exist for a matching that is a rational expectations solution: (i) no one matches in $t=1$, which would lead to $m_{\emptyset}$ (though note that $m_{\emptyset}$ is blocked by $(a_{1},b_{2})$ in $E_{\setminus\{a_{3},b_{1}\}}^{2}$), (ii) $a_{1}$ matches with $b_{2}$, and (iii) $a_{2}$ matches with $b_{2}$. Under (ii), note that the outcome of side-$A$ proposing deferred acceptance in $t=2$ among $a_{2},a_{3},a_{4},b_{1},b_{3},b_{4}$ is
$\left.\begin{array}[]{lcl}a_{2}&\_\_&b_{4}\\ a_{3}&\_\_&b_{3}\\ a_{4}&\_\_&b_{1}\end{array}\right.,$
so that any rational expectations solution for $E_{\setminus\{a_{3},b_{1}\}}^{2}$ that matches $a_{1},b_{2}$ in $t=1$ (if any) gives $b_{1}$ at least the payoff of matching with $a_{4}$. Similarly, under (iii), note that the outcome of side-$A$ proposing deferred acceptance in $t=2$ among $a_{1},a_{3},a_{4},b_{1},b_{3},b_{4}$ is
$\left.\begin{array}[]{lcl}a_{1}&\_\_&b_{4}\\ a_{3}&\_\_&b_{3}\\ a_{4}&\_\_&b_{1}\end{array}\right.,$
so that any rational expectations solution for $E_{\setminus\{a_{3},b_{1}\}}^{2}$ that matches $a_{2},b_{2}$ in $t=1$ (if any) gives $b_{1}$ at least the payoff of matching with $b_{4}$. It follows that in $E_{\setminus\{a_{3}\}}^{2}$, $b_{1}$ has to match with $a_{4}$.
Fourth, consider $E_{\setminus\{a_{2}\}}^{2}$. We argue that $m_{a_{2}}$ is a rational expectations solution for that economy. To see this, we need to argue that $b_{1}$ cannot block. Consider then the economy $E_{\setminus\{a_{2},b_{1}\}}^{2}$ and note that $m_{b_{1}}$ is a rational expectations solution for the same reason it is a rational expectations solution in $E_{\setminus\{b_{1}\}}^{2}$. Thus, by blocking $m_{a_{2}}$, $b_{1}$ cannot expect to get more than the payoff of matching with $a_{1}$ in $t=2$. It follows that $m_{a_{2}}$ is a rational expectations solution for $E_{\setminus\{a_{2}\}}^{2}$.
Finally, we argue that $m^{\star}$ in the example is the outcome of the procedure discussed in Section 4. Note that because of $m_{a_{2}}$, $a_{2}$ is willing to match with $b_{2}$ in the one-period economy induced by the conjectures, and similarly, $b_{1}$ is willing to match with $a_{1}$. Because $a_{2}$ and $b_{2}$ top-rank each other in the one-period economy induced by the conjectures, they match together in any stable matching, which then justifies $a_{1}$ and $b_{1}$ matching, and $a_{3}$ remaining unmatched. When $\{a_{3},a_{4}\}$ are the $t=2$-agents on side $A$, a unique stable matching exists, which is that specified by $m^{\star}_{2}$.
Appendix B Omitted proofs from Section 5
Proof of 5.1.
The proof of 5.1 proceeds by induction on the economy length, $T\in\mathbb{N}$. For $T=1$, the correspondence $\hat{\mathrm{DS}}_{1}$ is nonempty valued (cf. 1) and the statement about $\varphi_{\hat{\mathrm{DS}},1}$ is vacuous.
Assume then that we know the correspondence $\hat{\mathrm{DS}}_{t}$ is nonempty valued for $t\in\{1,\dots,T-1\}$. As it is clear from the main text, to establish the main result, showing that for any $E^{T}$, the period-1 conjectures $\{\varphi_{\hat{\mathrm{DS}},1}(\cdot,k):k\in A_{1}\cup B_{1}\}$ are nonempty and satisfy consistency is enough.
Fix an economy $E^{T}$. For each $k\in A_{1}\cup B_{1}$, let $\varphi_{0}(\emptyset,k)=\pazocal{M}_{T}(\emptyset,k)$ and recursively define for $n\geq 1$
$\varphi_{n}(\emptyset,k)=\left\{\overline{m}\in\pazocal{M}_{T}(\emptyset,k):\begin{array}[]{l}\overline{m}_{1}\in\mathrm{S}(E_{1,\varphi_{n-1}}^{\overline{m}_{1}}),(\overline{m}_{t})_{t=2}^{T}|_{E_{2}^{T}(\overline{m}_{1})}\in\hat{\mathrm{DS}}_{T-1}(E_{2}^{T}(\overline{m}_{1}))\end{array}\right\}.$
We observe that for all $k\in A_{1}\cup B_{1}$, the following hold. First, $\varphi_{n+1}(\emptyset,k)\subseteq\varphi_{n}(\emptyset,k)$ because in each step we are increasing the continuation values of period-1 agents other than $k$. Formally, by definition we have that for all $k\in A_{1}\cup B_{1}$, $\varphi_{1}(\emptyset,k)\subseteq\varphi_{0}(\emptyset,k)$. Suppose we have shown that $\varphi_{n}(\emptyset,k^{\prime})\subseteq\varphi_{n-1}(\emptyset,k^{\prime})$ for all $k^{\prime}\in A_{1}\cup B_{1}$; we proceed to show that $\varphi_{n+1}\subseteq\varphi_{n}$. Fix $k\in A_{1}\cup B_{1}$ and $\overline{m}\in\varphi_{n+1}(\emptyset,k)$. We only need to show that if $\overline{m}_{1}\in\mathrm{S}(E_{1,\varphi_{n}}^{\overline{m}_{1}})$, then $\overline{m}_{1}\in\mathrm{S}(E_{1,\varphi_{n-1}}^{\overline{m}_{1}})$. To this end, let $k^{\prime}$ be such that $\overline{m}_{1}(k^{\prime})\neq k^{\prime}$. By construction, $k^{\prime}$ prefers $\overline{m}$ to the worst conjectured matching in $\varphi_{n}(\emptyset,k^{\prime})\subseteq\varphi_{n-1}(\emptyset,k^{\prime})$. Thus, $k^{\prime}$ also prefers $\overline{m}$ to the worst conjectured matching in $\varphi_{n-1}(\emptyset,k^{\prime})$. Thus, $\overline{m}\in\varphi_{n}(\emptyset,k)$. Second, because everything is finite $\varphi_{\infty}(\emptyset,k)=\lim_{n\rightarrow\infty}\varphi_{n}(\emptyset,k)$ is well-defined. Third, because $\{\overline{m}\in\pazocal{M}_{T}(\emptyset,k):(\overline{m}_{t})_{t=2}^{T}\in\hat{\mathrm{DS}}_{T-1}\left(E_{2}^{T}(\overline{m}_{1})\right)\}\subseteq\varphi_{n}(\emptyset,k)$ for all $n$ and is nonempty, then $\varphi_{\infty}(\emptyset,k)$ is nonempty.
We claim that
$\varphi_{\infty}(\emptyset,k)=\left\{\overline{m}\in\pazocal{M}_{T}(\emptyset,k):\begin{array}[]{l}\overline{m}_{1}\in\mathrm{S}(E_{1,\varphi_{\infty}}^{\overline{m}_{1}}),(\overline{m}_{s})_{s=2}^{T}|_{E_{2}^{T}(\overline{m}_{1})}\in\hat{\mathrm{DS}}_{T-1}(E_{2}^{T}(\overline{m}_{1}))\end{array}\right\},$ (B.2)
and hence $\varphi_{\infty}=\varphi_{\hat{\mathrm{DS}},1}$.
To see that Equation B.2 holds, note the following:
  1. 1.
    If $\overline{N}$ exists such that for all $k\in A_{1}\cup B_{1}$, $\varphi_{\overline{N}}(\emptyset,k)=\varphi_{\overline{N}+1}(\emptyset,k)$, then for all $N\geq\overline{N}$ and all $k\in A_{1}\cup B_{1}$, it follows that $\varphi_{N}(\emptyset,k)=\varphi_{\overline{N}+1}(\emptyset,k)$. This follows from the definition of $\varphi_{n}$.
  2. 2.
    Because the sets $\varphi_{n}(\emptyset,k)$ are finite and so is the set of period-1 agents, $\overline{N}<\infty$ exists such that for all $k\in A_{1}\cup B_{1}$, $\varphi_{\overline{N}}=\varphi_{\overline{N}+1}$. Toward a contradiction, suppose that for all $N\in\mathbb{N}$ $k_{N}\in A_{1}\cup B_{1}$ exists such that $\varphi_{N}(\emptyset,k_{N})\neq\varphi_{N+1}(\emptyset,k_{N})$. Since the set $A_{1}\cup B_{1}$ is finite, we can find an agent $k$ and a subsequence $(n^{k})$ such that $\varphi_{n^{k}}(\emptyset,k)\neq\varphi_{n^{k}+1}(\emptyset,k)$. Thus, $\varphi_{n^{k}+1}(\emptyset,k)\subsetneq\varphi_{n^{k}}(\emptyset,k)$. Since $\varphi_{n^{k}}(\emptyset,k)$ is finite, this means that eventually $\varphi_{n^{k}}(\emptyset,k)=\emptyset$, a contradiction. From now on, let $\overline{N}$ denote the smallest $n$ such that for all $k\in A_{1}\cup B_{1}$, $\varphi_{\overline{N}}(\emptyset,k)=\varphi_{\overline{N}+1}(\emptyset,k)\equiv\varphi_{\infty}(\emptyset,k)$.
  3. 3.
    Suppose $k\in A_{1}\cup B_{1}$ exists such that $\overline{m}\in\varphi_{\infty}(\emptyset,k)$, but $\overline{m}$ is not an element of the set on the right-hand side of Equation B.2. Then, without loss of generality, $a\in A_{1}\cup B_{1}\setminus\{k\}$ exists such that $\overline{m}_{1}(a)\neq a$ and $\min_{\tilde{m}\in\varphi_{\infty}(\emptyset,a)}U_{1}(a,\tilde{m})>u(a,\overline{m}_{1}(a)).$ This implies that $\overline{m}\notin\varphi_{\overline{N}+2}(\emptyset,k)$, contradicting that $\varphi_{\overline{N}+1}(\emptyset,k)=\varphi_{\infty}(\emptyset,k)$.
We conclude that for all $k\in A_{1}\cup B_{1}$, Equation B.2 holds and hence $\varphi_{\infty}=\varphi_{\hat{\mathrm{DS}},1}$. Thus, for all $k\in A_{1}\cup B_{1}$, $\varphi_{\hat{\mathrm{DS}}}(\emptyset,k)\neq\emptyset$.
We now show that $\varphi_{\hat{\mathrm{DS}},1}$ satisfy CC. To this end, construct a candidate $m^{\star}$ as described in Section 4 using $\varphi_{\hat{\mathrm{DS}},1}$ and the correspondence $\hat{\mathrm{DS}}_{T-1}$, which by the inductive hypothesis is nonempty. Fix a period-1 agent, $k\in A_{1}\cup B_{1}$ such that $m^{\star}_{1}(k)=k$. Note that $m^{\star}$ satisfies the following: First, the continuation matching $(m^{\star}_{t})_{t=2}^{T}\in\hat{\mathrm{DS}}_{T-1}(E^{T}_{2}(m^{\star}_{1}))$. Second, for all $k^{\prime}\in A_{1}\cup B_{1}$ such that $m^{\star}_{1}(k^{\prime})\neq k^{\prime}$, we have that $k^{\prime}$ is matched to someone they prefer to the worst matching in $\varphi_{\hat{\mathrm{DS}},1}(\emptyset,k^{\prime})$. Finally, note that $m^{\star}_{1}$ has no blocking pairs $(a,b)$ such that $m^{\star}_{1}(a)\neq a,m^{\star}_{1}(b)\neq b$. Thus, $m^{\star}_{1}\in\mathrm{S}(E_{1,\varphi_{\hat{\mathrm{DS}}}}^{m^{\star}_{1}})$. Hence, $m^{\star}\in\varphi_{\hat{\mathrm{DS}},1}(\emptyset,k)$ so that CC holds.
Theorem 4.1 then implies that $m^{\star}\in\hat{\mathrm{DS}}_{T}(E^{T})$, which completes the inductive step. ∎